OpenStudy (anonymous):

Given the function f(x)=4(x+1)^2-3, indicate the shifts that will effects the location of the vertex and explain what effect they will have. Use complete sentences. 1. f(x-2) 2. f(x)-2 3. f(2x) 4. 2*f(x)

OpenStudy (anonymous):

@E.ali @phi

OpenStudy (phi):

First, notice that the vertex is at -1,-3 if you replace x with x-2 in your equation what do you get ?

OpenStudy (phi):

f(x)=4(x+1)^2-3 replace x with x-2. that means erase the x and put in x-2 in its place

OpenStudy (anonymous):

f(x-2)=4(x-2+1)^2-3 f(x-2)=4(x-1)^2-3 f(x-2)=4x-4^2-3 f(x-2)=4x-16-3 f(x-2)=4x-19 Is that correct? @phi Sorry about the late response. I was busy...